OpenStudy (anonymous):

Write the equation y - 2 = 4(x + 5) in standard form. A. 4x + y = 7 B. 4x - y = 3 C. 4x - y = -22

OpenStudy (anonymous):

y - 2 = 4x + 20 -22 = 4x - y :C

OpenStudy (anonymous):

4x - y = -22